A and B are ideal gases the molecular weight of a and B are in the ratio of 1:4 the pressure of a gas mixture containing equal weights of A and B is P atm what is the partial pressure in atm of B in the mixture 

Grade:12th pass

1 Answers

Arun
5 years ago
 
Dear student
 
Let Ma = a
Mb = b
 
Xb = m/4a /(m/4a + m/a) = 1/5
 
Pb = Xb * P total = P/5
 
Hope it helps
 
Regards
Arun (askIITians forum expert)